# Learning Mean Field Control on Sparse Graphs

### v1

@@ 5 Learning Algorithms
The two system approximation reduces the complexity of otherwise intractable large interacting systems on networks to the MFs of each degree. The system state at any time is then given by low-degree MFs $\mu^{1}_{t},\mu^{2}_{t},\ldots,\mu^{k^{*}}_{t}$ and high-degree MF $\mu^{\infty}_{t}$, briefly $\boldsymbol{\mu}_{t}\coloneqq(\mu^{1}_{t},\mu^{2}_{t},\ldots,\mu^{k^{*}}_{t},\mu^{\infty}_{t})$. Given a state $\boldsymbol{\mu}_{t}$, the possible state evolutions depend only on the analogous set of low-degree and high-degree policies at that time, $\boldsymbol{\pi}_{t}\coloneqq(\pi^{1}_{t},\pi^{2}_{t},\ldots,\pi^{k^{*}}_{t},\pi^{\infty}_{t})$. Therefore, choosing a $\boldsymbol{\pi}_{t}$ fully defines the state transition of the overall system, and is thus considered as the high-level action in the MFC MDP. Introducing a high-level policy $\hat{\pi}$ to output $\boldsymbol{\pi}_{t}\sim\hat{\pi}_{t}(\boldsymbol{\pi}_{t}\mid\boldsymbol{\mu}_{t})$ allows us to solve for an optimal set of policies by solving the MFC MDP for optimal $\hat{\pi}$, since the limiting MF dynamics are deterministic. Finally, the MFC MDP is solved by applying single-agent policy gradient RL, resulting in Algorithm 1. In practice, we use proximal policy optimization (Schulman et al., 2017). To lower the complexity of the resulting MDP, we parametrize policies as distributions over actions given the node state, $\pi^{k}_{t}\in\mathcal{P}(\mathcal{U})^{\mathcal{X}}$.
@@
[…]
@@ 6 Examples
For a general overview of many applications, see Laurière et al. (2022a). We consider four problems briefly described here. Problem details can be found in Appendix C.
@@
The classical SIS model (Kermack & McKendrick, 1927; Brauer, 2005) is a benchmark in the MF learning literature (Laurière et al., 2022b; Zhou et al., 2024). Agents are infected or susceptible, resulting in the state space $\mathcal{X}\coloneqq\{S,I\}$, and decide to protect themselves or not. The infection probability increases without protection, and with the number of infected neighbors. Furthermore, there is a constant probability for recovering from an infection. The SIR model (Hethcote, 2000; Doncel et al., 2022) is an extension of SIS where agents can also be in a recovered state $R$ where they are immune to a reinfection. Consequently, the state space for the SIR model is $\mathcal{X}\coloneqq\{S,I,R\}$.
@@
Inspired by graph coloring (Jensen & Toft, 2011; Barenboim & Elkin, 2022), the states are finitely many colors on a circle and a target color distribution is given. Agents stay at their color or costly move to a neighboring color. The objective decreases for deviations from the target color distribution and if neighbors of an agent have neighboring colors to the agent’s color on the respective color ring.
@@
In the rumor model (Maki & Thompson, 1973; Gani, 2000; Cui et al., 2022), agents are either aware of a rumor and are consequently in the aware state $A$ or they have not heard the rumor and are in the ignorant state $I$. Aware agents decide whether they spread the rumor to their neighbors or keep it to themselves. They are awarded for spreading the rumor to unaware agents but loose reputation for telling the rumor to already aware agents.
Table 2: (LW)MFC policy gradient, (LW)MFMARL policy gradient, and IPPO for four problems on synthetic CL graphs of size $N$. Best objective after 24 hours of training on 96 CPUs. Best result for each problem-graph tuple in bold.
@@ 7 Simulation & Results
In this section, we numerically verify the two system approximation as well as the proposed learning algorithms by comparing them with baselines from the literature. The two systems approximation is compared with previous graph approximations such as graphex or Lp graphon MF equations, and the learning algorithms are verified against standard scalable independent learning methods such as IPPO (Tan, 1993; Papoudakis et al., 2021), due to the large scale of networks considered here. To generate artificial networks of different sizes we employ a CL-based graph sampling algorithm (Chung & Lu, 2002; Miller & Hagberg, 2011) from the Python NetworkX package.
We compare the accuracy of our model on different empirical datasets with Lp graphon and graphex based models and with our extensive approximation LWMFC, where computationally feasible, to see how much information is lost in the LWMFC approximation. We use eight datasets from the KONECT database (Kunegis, 2013), where we substitute directed or weighted edges by simple undirected edges: CAIDA (Leskovec et al., 2007)($N\approx 26k$), Cities (Kunegis, 2013) ($N\approx 14k$), Digg Friends (Hogg & Lerman, 2012) ($N\approx 280k$), Enron (Klimt & Yang, 2004) ($N\approx 87k$), Flixster (Zafarani & Liu, 2009) ($N\approx 2.5mm$), Slashdot (Gómez et al., 2008) ($N\approx 50k$), Yahoo (Kunegis, 2013) ($N\approx 653k$), and YouTube (Mislove, 2009) ($N\approx 3.2mm$). See the references for details on the respective empirical networks.
@@ Results.
First, we establish the usefulness of LWMFC and LWMFC by comparing their dynamics to those of LPGMFGs and GXMFGs (Fabian et al., 2023, 2024) on eight real-world networks, see Figure 2 for examplary dynamics over time. As Table 1 shows, our LWMFC approach clearly outperforms the current LPGMFGs and GXMFGs methods for all empirical networks and problems. The extensive approximation LWMFC moderately outperforms LWMFC across datasets, except Yahoo. Since the extensive approximation is more detailed, it is often more accurate than the LWMFC approximation. However, Table 1 lacks an evaluation for LWMFC on the Color and Rumor problem because the extensive approximation is computationally too expensive for these problems. Consequently, LWMFC dynamics are the more practical choice since they are computationally tractable and yield a good performance across problems and datasets.
@@ Appendix A Appendix: Proofs for the Theoretical Results
We aim to eventually apply Lacker et al. (2023, Theorem 3.6) and therefore have to check that the respective conditions hold in our model. Keeping in mind Assumption 2.2 and the i.i.d. initial distribution $\mu_{0}$, we leverage Lacker et al. (2023, Corollary 2.16) to obtain convergence in probability in the local weak sense of the marked graphs $(G_{N},X^{G_{N}})$ to the limiting marked graph $(G,X^{G})$.
Since the theory in Lacker et al. (2023) is only formulated in terms of particle systems without including actions in the form of policies, we provide a suitable reformulation of our cooperative mean field game model. Thus, define an auxiliary extended state space $\mathcal{X}_{\mathrm{e}}\coloneqq\mathcal{X}\cup(\mathcal{X}\times\mathcal{U})$ which serves as the state space for the extended particle system for some fixed policy ensemble $\pi$. The idea behind the extended state space $\mathcal{X}_{\mathrm{e}}$ is to define an extended particle system where the state transition in $\mathcal{X}$ and the choice of the next action $u_{t+1}\in\mathcal{U}$ are separated into two different time steps.
Using the notations from Lacker et al. (2023), denote by $\mathcal{S}^{\sqcup}(\mathcal{X})$ the set of finite unordered sequences of arbitrary length with values in $\mathcal{X}$ and by $\Xi\coloneqq\mathcal{X}^{\mathcal{X}\times\mathcal{S}^{\sqcup}(\mathcal{X})}\times\mathcal{U}^{\mathcal{X}\times\mathcal{U}\times\mathcal{S}^{\sqcup}(\mathcal{X})}$ the set of possible noise values. Next, specify a transition function $F^{\tau}:\mathcal{X}_{\mathrm{e}}\times\mathcal{S}^{\sqcup}(\mathcal{X}_{\mathrm{e}})\times\Xi\to\mathcal{X}_{\mathrm{e}}$ for each $\tau\in\mathcal{T}_{\mathrm{e}}\coloneqq\{0\}\cup[2T-1]$ by
@@
Now, it remains to check that Lacker et al. (2023, Assumption A) is satisfied by the extended particle system defined above. First, the noise terms $\xi_{i,\tau}$ are i.i.d. distributed for all agents $i\in[N]$ and with respect to all time points $\tau\in\mathcal{T}_{\mathrm{e}}$ by construction. Finally, keeping in mind that the respective spaces are discrete, the map $F^{\tau}$ is continuous for each $\tau\in\mathcal{T}_{\mathrm{e}}$. Therefore, Lacker et al. (2023, Theorem 3.6) yields the desired result. ∎
@@
The reward $r$ is a continuous function by Assumption 3.2. Furthermore, by Theorem 3.1 we know that the empirical mean fields converge in probability to the limiting mean fields. Hence, we can apply the continuous mapping theorem (Mann & Wald, 1943; Van der Vaart, 2000) to obtain the desired result. ∎
@@
Since we want to lower bound the number of possible $t$-hop neighborhoods $N_{G,t}$, we assume for simplicity that $t$-hop neighbors of the initial agent have at most degree $k$ themselves. Furthermore, we keep in mind the fact (Beck & Robins, 2007, Theorem 2.2) that in a $d$-dimensional simplex with edge length $\ell\in\mathbb{N}$, the number of integer points contained in the simplex is
@@
Since Lemma 4.1 considers the worst case, it suffices to prove the lower bound $\Omega\left(2^{\mathrm{poly}(k)}\right)$ for one class of CL graphs. We choose our running example of CL power law graphs with coefficient above two. It is well known that these large power law CL graphs are locally tree-like (Van Der Hofstad, 2024, Theorem 3.18) which we tacitly exploit in the following induction proof.
The proof is via induction over $t$. We start with $t=1$ and the corresponding $1$-hop neighborhood. The neighborhood consists of $k$ agents where each one has a degree in $[k]$ and a state in $\mathcal{X}$. Since the agents themselves are indistinguishable in our model, we focus on the degree-state neighborhood distributions. Then, the set of possible state-degree neighborhood distributions can be seen as the integer points in a $(k+|\mathcal{X}|-1)$-dimensional simplex with edge length $k$. Keeping in mind Equation (1) and the well-known Stirling approximation, see e.g. Marsaglia & Marsaglia (1990), we obtain
@@
Now, it remains to establish the induction step from $t$ to $t+1$ where we assume that $N_{G,t}\in\Omega\left(2^{\mathrm{poly}(k)}\right)$ holds. Then, instead of looking at the $(t+1)$-hop neighborhoods of the initial agent, we can equivalently look at his or her $1$-hop neighborhoods where each neighbor’s ’extended state’ now consists of the neighbor’s $t$-hop neighborhood, where we ignore the edge between the neighbor and initial agent. Thus, the simplex edge length decreases by one from $k$ to $k-1$ which is negligible for large $k$. Leveraging the induction assumption, we obtain
@@
=$\overset{\mathrm{Stirling}}{=}\Omega\left(\sqrt{\frac{N_{G,t}+k}{kN_{G,t}}}\frac{(N_{G,t}+k)^{N_{G,t}+k}}{N_{G,t}^{N_{G,t}}k^{k}}\right)$
@@
=$\overset{\mathrm{(IA)}}{=}\Omega\left(\frac{1}{\sqrt{k}}\frac{\left(2^{\mathrm{poly}(k)}\right)^{k}}{k^{k}}\right)$
@@ Appendix B Extensive Approximation Derivation
(2)$\ \approx\end{subarray}\frac{P\left(\deg(v_{1})=m\mid(v_{0},v_{1})\in E\right)\mu_{t}^{m}(s)}{P\left(\deg(v_{1})>k^{*}\cap x_{t}^{1}=s\mid(v_{0},v_{1})\in E\right)+\sum_{k=1}^{k^{*}}P\left(\deg(v_{1})=k\cap x_{t}^{1}=s\mid(v_{0},v_{1})\in E\right)}$
@@
(2)$\ \approx\end{subarray}P\left(\deg(v_{1})>k^{*}\mid(v_{0},v_{1})\in E\right)\mu_{t}^{\infty}(s)+\sum_{k=1}^{k^{*}}P\left(\deg(v_{1})=k\mid(v_{0},v_{1})\in E\right)\mu_{t}^{k}(s)\,.$ ) .
@@
[…]
@@
[…]
@@
${}\displaystyle\quad\cdot\sum_{c\in\boldsymbol{\mathcal{C}}^{k}}\left[\sum_{\boldsymbol{a}_{2}\in\boldsymbol{\mathcal{A}}^{k}(G^{\prime},c)}\prod_{j=1}^{d}\binom{g_{j}^{\prime}}{a_{j1},\ldots,a_{j\infty}}\prod_{m\in[k^{*}]\cup\{\infty\}}\right.$

### v2

@@ 5 Learning Algorithms
The two system approximation reduces the complexity of otherwise intractable large interacting systems on networks to the MFs of each degree. The system state at any time is then given by low-degree MFs $\mu^{1}_{t},\mu^{2}_{t},\ldots,\mu^{k^{*}}_{t}$ and high-degree MF $\mu^{\infty}_{t}$, briefly $\boldsymbol{\mu}_{t}\coloneqq(\mu^{1}_{t},\mu^{2}_{t},\ldots,\mu^{k^{*}}_{t},\mu^{\infty}_{t})$. Given a state $\boldsymbol{\mu}_{t}$, the possible state evolutions depend only on the analogous set of low-degree and high-degree policies at that time, $\boldsymbol{\pi}_{t}\coloneqq(\pi^{1}_{t},\pi^{2}_{t},\ldots,\pi^{k^{*}}_{t},\pi^{\infty}_{t})$. Therefore, choosing a $\boldsymbol{\pi}_{t}$ fully defines the state transition of the overall system, and is thus considered as the high-level action in the MFC MDP. Introducing a high-level policy $\hat{\pi}$ to output $\boldsymbol{\pi}_{t}\sim\hat{\pi}_{t}(\boldsymbol{\pi}_{t}\mid\boldsymbol{\mu}_{t})$ allows us to solve for an optimal set of policies by solving the MFC MDP for optimal $\hat{\pi}$, since the limiting MF dynamics are deterministic. Finally, the MFC MDP is solved by applying single-agent policy gradient RL, resulting in Algorithm 1. In practice, we use proximal policy optimization (schulman2017proximal). To lower the complexity of the resulting MDP, we parametrize policies as distributions over actions given the node state, $\pi^{k}_{t}\in\mathcal{P}(\mathcal{U})^{\mathcal{X}}$.
@@
Model Average expected total variation $\Delta\mu$ in $\%$, standard deviation in brackets
@@ 6 Examples
For a general overview of many applications, see lauriere2022learning. We consider four problems briefly described here. Problem details can be found in Appendix C.
@@
The classical SIS model (kermack1927contribution; brauer2005kermack) is a benchmark in the MF learning literature (lauriere2022scalable; zhou2024graphon). Agents are infected or susceptible, resulting in the state space $\mathcal{X}\coloneqq\{S,I\}$, and decide to protect themselves or not. The infection probability increases without protection, and with the number of infected neighbors. Furthermore, there is a constant probability for recovering from an infection. The SIR model (hethcote2000mathematics; doncel2022mean) is an extension of SIS where agents can also be in a recovered state $R$ where they are immune to a reinfection. Consequently, the state space for the SIR model is $\mathcal{X}\coloneqq\{S,I,R\}$.
@@
Inspired by graph coloring (jensen2011graph; barenboim2022distributed), the states are finitely many colors on a circle and a target color distribution is given. Agents stay at their color or costly move to a neighboring color. The objective decreases for deviations from the target color distribution and if neighbors of an agent have neighboring colors to the agent’s color on the respective color ring.
@@
In the rumor model (maki1973mathematical; gani2000maki; cui2022hypergraphon), agents are either aware of a rumor and are consequently in the aware state $A$ or they have not heard the rumor and are in the ignorant state $I$. Aware agents decide whether they spread the rumor to their neighbors or keep it to themselves. They are awarded for spreading the rumor to unaware agents but loose reputation for telling the rumor to already aware agents.
Table 2: (LW)MFC policy gradient, (LW)MFMARL policy gradient, and IPPO for four problems on synthetic CL graphs of size NN. Best objective after 24 hours of training on 96 CPUs. Best result for each problem-graph tuple in bold.
@@ 7 Simulation & Results
In this section, we numerically verify the two system approximation as well as the proposed learning algorithms by comparing them with baselines from the literature. The two systems approximation is compared with previous graph approximations such as graphex or Lp graphon MF equations, and the learning algorithms are verified against standard scalable independent learning methods such as IPPO (tan1993multi; papoudakis1benchmarking), due to the large scale of networks considered here. To generate artificial networks of different sizes we employ a CL-based graph sampling algorithm (chung2002connected; miller2011efficient) from the Python NetworkX package.
We compare the accuracy of our model on different empirical datasets with Lp graphon and graphex based models and with our extensive approximation LWMFC, where computationally feasible, to see how much information is lost in the LWMFC approximation. We use eight datasets from the KONECT database (kunegis2013konect), where we substitute directed or weighted edges by simple undirected edges: CAIDA (leskovec2007graph)($N\approx 26k$), Cities (kunegis2013konect) ($N\approx 14k$), Digg Friends (hogg2012social) ($N\approx 280k$), Enron (klimt2004enron) ($N\approx 87k$), Flixster (Zafarani+Liu:2009) ($N\approx 2.5mm$), Slashdot (gomez2008statistical) ($N\approx 50k$), Yahoo (kunegis2013konect) ($N\approx 653k$), and YouTube (mislove2009online) ($N\approx 3.2mm$). See the references for details on the respective empirical networks.
@@ Results.
First, we establish the usefulness of LWMFC and LWMFC by comparing their dynamics to those of LPGMFGs and GXMFGs (fabian2023learning; fabian2024learning) on eight real-world networks, see Figure 2 for examplary dynamics over time. As Table 1 shows, our LWMFC approach clearly outperforms the current LPGMFGs and GXMFGs methods for all empirical networks and problems. The extensive approximation LWMFC moderately outperforms LWMFC across datasets, except Yahoo. Since the extensive approximation is more detailed, it is often more accurate than the LWMFC approximation. However, Table 1 lacks an evaluation for LWMFC on the Color and Rumor problem because the extensive approximation is computationally too expensive for these problems. Consequently, LWMFC dynamics are the more practical choice since they are computationally tractable and yield a good performance across problems and datasets.
@@ Appendix A Appendix: Proofs for the Theoretical Results
We aim to eventually apply lacker2023local and therefore have to check that the respective conditions hold in our model. Keeping in mind Assumption 2.2 and the i.i.d. initial distribution $\mu_{0}$, we leverage lacker2023local to obtain convergence in probability in the local weak sense of the marked graphs $(G_{N},X^{G_{N}})$ to the limiting marked graph $(G,X^{G})$.
Since the theory in lacker2023local is only formulated in terms of particle systems without including actions in the form of policies, we provide a suitable reformulation of our LWMFC model. Thus, define an auxiliary extended state space $\mathcal{X}_{\mathrm{e}}\coloneqq\mathcal{X}\cup(\mathcal{X}\times\mathcal{U})$ which serves as the state space for the extended particle system for some fixed policy ensemble $\pi$. The idea behind the extended state space $\mathcal{X}_{\mathrm{e}}$ is to define an extended particle system where the state transition in $\mathcal{X}$ and the choice of the next action $u_{t+1}\in\mathcal{U}$ are separated into two different time steps.
Using the notations from lacker2023local, denote by $\mathcal{S}^{\sqcup}(\mathcal{X})$ the set of finite unordered sequences of arbitrary length with values in $\mathcal{X}$ and by $\Xi\coloneqq\mathcal{X}^{\mathcal{X}\times\mathcal{S}^{\sqcup}(\mathcal{X})}\times\mathcal{U}^{\mathcal{X}\times\mathcal{U}\times\mathcal{S}^{\sqcup}(\mathcal{X})}$ the set of possible noise values. Next, specify a transition function $F^{\tau}:\mathcal{X}_{\mathrm{e}}\times\mathcal{S}^{\sqcup}(\mathcal{X}_{\mathrm{e}})\times\Xi\to\mathcal{X}_{\mathrm{e}}$ for each $\tau\in\mathcal{T}_{\mathrm{e}}\coloneqq\{0\}\cup[2T-1]$ by
@@
Now, it remains to check that lacker2023local is satisfied by the extended particle system defined above. First, the noise terms $\xi_{i,\tau}$ are i.i.d. distributed for all agents $i\in[N]$ and with respect to all time points $\tau\in\mathcal{T}_{\mathrm{e}}$ by construction. Finally, keeping in mind that the respective spaces are discrete, the map $F^{\tau}$ is continuous for each $\tau\in\mathcal{T}_{\mathrm{e}}$. Therefore, lacker2023local yields the desired result. ∎
@@
The reward $r$ is a continuous function by Assumption 3.2. Furthermore, by Theorem 3.1 we know that the empirical mean fields converge in probability to the limiting mean fields. Hence, we can apply the continuous mapping theorem (mann1943stochastic; van2000asymptotic) to obtain the desired result. ∎
@@
Since we want to lower bound the number of possible $t$-hop neighborhoods $N_{G,t}$, we assume for simplicity that $t$-hop neighbors of the initial agent have at most degree $k$ themselves. Furthermore, we keep in mind the fact (beck2007computing, Theorem 2.2) that in a $d$-dimensional simplex with edge length $\ell\in\mathbb{N}$, the number of integer points contained in the simplex is
@@
Since Lemma 4.1 considers the worst case, it suffices to prove the lower bound $\Omega\left(2^{\mathrm{poly}(k)}\right)$ for one class of CL graphs. We choose our running example of CL power law graphs with coefficient above two. It is well known that these large power law CL graphs are locally tree-like (van2024random, Theorem 3.18) which we tacitly exploit in the following induction proof.
The proof is via induction over $t$. We start with $t=1$ and the corresponding $1$-hop neighborhood. The neighborhood consists of $k$ agents where each one has a degree in $[k]$ and a state in $\mathcal{X}$. Since the agents themselves are indistinguishable in our model, we focus on the degree-state neighborhood distributions. Then, the set of possible state-degree neighborhood distributions can be seen as the integer points in a $(k+|\mathcal{X}|-1)$-dimensional simplex with edge length $k$. Keeping in mind Equation (1) and the well-known Stirling approximation, see e.g. marsaglia1990new, we obtain
@@
Now, it remains to establish the induction step from $t$ to $t+1$ where we assume that $N_{G,t}\in\Omega\left(2^{\mathrm{poly}(k)}\right)$ holds. Then, instead of looking at the $(t+1)$-hop neighborhoods of the initial agent, we can equivalently look at his or her $1$-hop neighborhoods where each neighbor’s ’extended state’ now consists of the neighbor’s $t$-hop neighborhood, where we ignore the edge between the neighbor and initial agent. Thus, the simplex edge length decreases by one from $k$ to k1k-1 which is negligible for large $k$. Leveraging the induction assumption, we obtain
@@
$\overset{\mathrm{Stirling}}{=}\Omega\left(\sqrt{\frac{N_{G,t}+k}{kN_{G,t}}}\frac{(N_{G,t}+k)^{N_{G,t}+k}}{N_{G,t}^{N_{G,t}}k^{k}}\right)$
@@
$\overset{\mathrm{(IA)}}{=}\Omega\left(\frac{1}{\sqrt{k}}\frac{\left(2^{\mathrm{poly}(k)}\right)^{k}}{k^{k}}\right)$
@@ Appendix B Extensive Approximation Derivation
[…]
@@
[…]
@@
$\qquad\qquad\qquad\left.\sum_{m^{\prime}\in[k^{*}]\cup\{\infty\}}a_{jm^{\prime}}=g^{\prime}_{j},\forall j\in[d]\quad\textrm{and}\quad\sum_{\ell=1}^{d}a_{\ell m}=c_{m},\forall m\in[k^{*}]\cup\{\infty\}\right\}\,.$
@@
 $\left.\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\textrm{ and }\sum_{\ell,\ell^{\prime}=1}^{d}a_{\ell\ell^{\prime}m}=c_{m},\forall m\in[k^{*}]\cup\{\infty\}\right\}\,.$
@@
$\quad\cdot\sum_{c\in\boldsymbol{\mathcal{C}}^{k}}\left[\sum_{\boldsymbol{a}_{2}\in\boldsymbol{\mathcal{A}}^{k}(G^{\prime},c)}\prod_{j=1}^{d}\binom{g_{j}^{\prime}}{a_{j1},\ldots,a_{j\infty}}\prod_{m\in[k^{*}]\cup\{\infty\}}\right$.